# ReasonVQA: A Multi-hop Reasoning Benchmark with Structural Knowledge for Visual Question Answering

### v1

@@ 3 The ReasonVQA Dataset
The ReasonVQA dataset is created to assess the capabilities of VQA models to answer multi-hop questions and to utilize external knowledge. It contains a large number of questions categorized into three levels of complexity: 1-hop, 2-hop, and 3-hop. To answer these questions, models must possess a robust ability to reason across multiple external facts. The overview of our framework, illustrated in Figure 2, to create ReasonVQA consists of three steps: (1) External Knowledge Integration, (2) Question Generation, (3) Dataset Construction. In the External Knowledge Integration step (Sect.3.1) we “link” annotated objects in the images to corresponding entities in external knowledge graphs. In the Question Generation step, multiple-hop questions are generated semi-automatically by traveling through the knowledge graph (Sect.3.2). In the Dataset Construction step (Sect.3.3), the question distribution is balanced to reduce bias, and all the images and questions are split into train set and test set with a similar answer distribution. The detailed workflow are described as algorithms in the supplementary materials.
@@ 3.2.1 Template-based Question Generation
Figure 3 shows the overview of our question generation process. In our template-based approach, questions are generated using predefined templates containing one or multiple placeholders to fill in. A template can generate many questions, helping to build low-cost and scalable systems. Although the templates are manually created, they are designed to be generic enough to provide extensive coverage. Our proposed method has the capability to combine multiple templates to generate a single question, enhancing the scalability of our system. Furthermore, as we construct potential templates based on knowledge gleaned from the knowledge base (KB), the diversity of the templates aligns with the breadth of knowledge available in the KB.
One-hop Question Generation. Let us assume we are constructing questions around an object in an image, referred to as the main object. After obtaining external facts from Wikidata for the main object as described in Section 3.1, we use a predefined template for each property (or relation, interchangeably) occuring in these facts. Each template contains a placeholder, which will be filled later by the class name of the main object, resulting in 1-hop question. For Wikidata, the class name of an object is simply the value of either the “instance of” or “subclass of” property. The Figure 4 shows an example of 1-hop template question. In this case, the template is “Who designed __?” and the class name “skyscraper” is used to fill in the placeholder, producing the complete question “Who designed this skyscraper?”. Note that during the process of retrieving facts about the main object from Wikidata, we also obtain the answer to the question. For example, the answer for the mentioned question is “César Pelli”.
@@ 3.2.2 False Choice Generation
To generate challenging multiple choice questions, in addition to the correct answer, we need to create non-trivial false choices. First, we assign a question into one of four categories based on its correct answer: fixed, date, number, and literal. The fixed category contains questions that have an enclosed set of choices, such as questions about gender, continent, country, and language. The date, number and literal categories contain questions about date time, numeric values, and literal values, respectively. For the questions in the fixed categories, the false choices are randomly selected from the corresponding set of choices. For the date category, the false choices are generated by randomly selecting $N$ date values within $[-10,+10]$ years. For the number category, the false choices are generated by randomly selecting $N$ numbers within a specific range $[m,n]$, where $m=\frac{i}{2},$ $n=max(1.5i,$ $m+2N)$ with $i$ being the correct answer. For the literal category, the false choice is generated by retrieving $N$ values that come from the same property as the correct answer. In practice, we set $N=3$, meaning that there is up to three false choices for each correct answer.
@@ 4.1 Dataset Statistics and Quality Evaluation
The latest version of ReasonVQA consists of more than 598K images and about 4.2M questions. The scalability of our dataset is based on two facts: (1) many different objects share the same attribute and (2) multi-hop questions are built by nesting multiple templates. For example, facts about buildings often contain the attribute “height” and we only need to define a template for this property one time during the generation process. To assess state-of-the-art approaches, we derived a subset version of our dataset, ReasonVQA-U. The main objective is to create a manageable subset of our dataset to assess and compare these models efficiently. The answer distribution balancing process was applied on this subset to obtain the balanced version ReasonVQA-B.
@@
In Table 1, we compare our dataset with various VQA datasets. In terms of the number of images, ours has the highest quantity. Moreover, we have a significantly larger number of questions compared to all the others, except GQA. To assess the quality of our dataset, we conducted a user study over 1000 randomly selected question and image pairs. The study involves 20 participants, who use English as the main working and/or studying language, to evaluate the natural level of 50 randomly selected questions each. They were asked to rate the naturalness on a four-level scale: (1) very unnatural; (2) unnatural; (3) natural; and (4) very natural. Additionally, participants also marked any questions with grammatical errors. The findings showed that 2.20% of the questions were rated as “very unnatural,” 13.9% as “unnatural,” 58.1% as “natural,” and 25.8% as “very natural,” with only 2.5% containing grammatical errors. This user study indicates that more than 83% of the questions are natural to humans, demonstrating that our dataset already has a good level of naturalness but still has room for improvement. Additional dataset statistics and results of the user study can be found in the Section 2 of the supplementary materials.
@@ 4.2 Baseline Experiments and Analysis
Fine-tuning and Performance Gains. To assess the effectiveness of fine-tuning on the dataset, we fine-tune three vision-language models of varying scales and compare their performance before and after training. This experiment evaluates how well models can adapt to the dataset and improve their understanding of knowledge-based visual question answering. By analyzing performance gains, we demonstrate the potential of fine-tuning in enhancing model capabilities beyond their zero-shot performance.
Impact of Dataset Scale on Model Performance. To assess the performance and scalability of LLMs, we also benchmark them using the ReasonVQA at multiple size increments. By gradually increasing the number of questions and images, we aim to observe how each model’s accuracy relative to the amount of data. This approach will allow us to identify any limitations in model capacity, potential bottlenecks, and the overall robustness of each model under different data volumes.
@@ 4.3 Metrics
We report the accuracy of the LLMs to assess their performance. In open-ended scenario, where the LLM outputs are in free-text format, comparing responses directly with the ground truth answers can be challenging. To address this, we tested several string-matching methods to compare the output with the ground truth answer. We employ three approaches - exact match, substring, and semantic similarity - which offer different strategies for aligning and scoring the outputs based on varying degrees of match tolerance and answer format flexibility. The exact match score is defined for each question as 1 if the answer provided by a mode is exactly the same as the ground truth answer; the score is 00 otherwise. However, foundation-based models usually provide answers in a form of free text. Therefore, the exact score might not correctly reflects the performance of the models. The substring score is defined as 1 if the answer provided by the model is contained in the list of words from the ground truth answer, and 00 otherwise. However, this approach falls short in many cases; for example, when the ground truth is “male” and the prediction is “a man,” the substring score is 00, which is incorrect. Therefore, the semantic similarity score is necessary to capture the semantic quality of the generated answer and provide a more meaningful evaluation of the LLMs. The score is based on the similarity of the generated answer and the ground truth computed using a widely recognized similarity all-MiniLM-L6-v2 [26], which is a compact, efficient model for semantic textual similarity tasks. The reported scores are the average values over the entire dataset. Section 3 of the supplementary material provides a more comprehensive report of the inference results.
@@ 4.4 Experimental Results
For recent models, we also reported in Figure 5 the accuracies in four string-matching approaches across different categories: the number of hop in the questions (1-hop, 2-hop, 3-hop); the availability of scene graph and the image source. Our results revealed that the overall accuracies computed by exact matching are quite low because these models treat each question as open-ended and respond in free text, leading to varying answers that need careful comparison with the ground truth. Their responses are often not entirely incorrect but lack detail. For example, instead of saying “390 meters” for the height of a tower, a model might say “very high,” which is not specific enough. The substring approach, which considers all words from both the predicted answers and the ground truth, yields more positive scores. Scores calculated using the semantic similarity matching are consistently higher because this method considers synonymous expressions, offering a more adaptable evaluation that closely aligns with human judgment.
@@
Figure 6: Performance of models across different dataset sizes in open-ended (top) and multiple choice (bottom) scenarios, with accuracy reported.
@@
Figure 6 illustrates the performance of various models across different dataset sizes. We observe that as the dataset grows, model performance initially improves but then declines with varying degrees of intensity. This decline likely occurs because as new samples are introduced, they bring additional domain knowledge and more complex questions, which may challenge the ability of models to maintain efficiency and accuracy under the increased data load. This observation suggests that integrating a wider variety of image sources and knowledge domains could create a more demanding benchmark, providing a more comprehensive assessment of model robustness.
@@ 5 Conclusions
We have proposed a novel VQA dataset in which external knowledge is required to answer questions. Our dataset construction framework is cost-effective, scalable, and requires minimal manual effort. To build this dataset, we leveraged existing annotations from CV datasets and external knowledge base Wikidata. Although our current dataset was built from specific image sources, our framework is generic enough to be expanded to integrate images from other CV datasets and maintain the diversity of the questions. We also offer a feature that allows users to select from 20 predefined domains, encompassing a broad range of human knowledge, to tailor the generated dataset to specific areas of interest. This mechanism enables users to focus the dataset on particular domains relevant to their needs. ReasonVQA covers a wide variety of knowledge and consists of complex questions assembled by joining the scene graph and multi-hop knowledge graph. Our dataset has been proven to be very challenging to state-of-the-art VQA models including those built with LLMs. We have also provided source code that enables easy integration of new images and knowledge sources for generating multi-hop questions. This allows researchers to expand and customize their own datasets. Thus, we believe that it will be a valuable resource to evaluate existing models and encourage the research community to develop more robust and powerful VQA systems.
6 Acknowledgements
@@ 1.1 Template-based Question Generation
2$\{\mathcal{C}_{i}\}$ = set of Wikidata entities corresponding to annotated objects in Img
@@ 12             $t\leftarrow\mathcal{T}_{s}[e_{j}]$
14      return $t\cup\leavevmode\nobreak\ \textnormal{{Generate(}}\textnormal{\emph{$e_{j+1}$}}\textnormal{{)}}$
@@ 19       $\{e_{j}\}$ = set of edges from $v$ to $\mathcal{C}_{i}$
20      $\mathcal{D}_{1}$ = $\mathcal{D}_{1}$cup  Generate($e_{0}$)
@@ 2.2 Dataset Domains
We categorize each question into one or more domains based on the property used to generate the question. For instance, a question about the birthday of a famous person, derived from the property “date of birth,” can be classified under both the “person” and “history-events” domains. This enables the system to receive input as one or multiple specified domains and generate questions accordingly. This mechanism is particularly useful for narrowing down the dataset to a specific domain, facilitating more focused and refined results. To achieve this, we predefined 20 domains covering all aspects of human knowledge, as outlined in the list below. During the question generation process, we select only the properties associated with the queried domains and filter out others, thereby optimizing the efficiency of question generation. Figure 7 presents the distribution of questions across domains in ReasonVQA.
@@ 2.3 Question Evaluation by User Study
To evaluate the quality of generated questions, we conducted a user study with 1,000 randomly selected question and image pairs. Twenty participants, all proficient in English as their primary language for work or study, assessed the naturalness of 50 randomly chosen questions each. They rated the questions on a four-level scale: (1) very unnatural, (2) unnatural, (3) natural, and (4) very natural. Additionally, they were also asked to mark any questions with grammatical errors. The results indicated that 2.2% of the questions were rated as “very unnatural,” 13.9% as “unnatural,” 58.1% as “natural,” and 25.8% as “very natural,”, with only 2.5% of the questions had grammatical errors. The results are visualized in Figure 8.
@@ Appendix 3 Experiments
In this section, we provide additional metrics, including the Standard Deviation (SD) and the Standard Error of the Mean (SEM), for the benchmarked models. These scores are further analyzed across various aspects of our dataset.
3.1 Experiment Results
We evaluated the performance of 13 large language models (LLMs) BLIP-2 [18], InstructBLIP [5], mPLUG-Owl2 [36], Idefics2 [16], Mantis-SigLIP [13], Mantis-Idefics2 [13], mPLUG-Owl3 [35], GPT-4o [12], LLaVA-OV [17], Qwen2.5-VL [2], PaliGemma-2 [30], PaliGemma-2-Mix [24], SmolVLM-Instruct [21] using the SD and SEM scores, as presented in Tables 3 and 4. Overall, the SD values are relatively large, indicating a high degree of variability in the model performances across different aspects of our dataset. This variability indicates that ReasonVQA poses significant challenges to the models, as large SD values typically reflect a high sensitivity to the diverse types of samples or tasks within the dataset. At the same time, the relatively low SEM values (ranging from 0.1% to 0.7%) suggest that the mean accuracy scores for each model are estimated with high precision, despite the substantial variability indicated by the SD. This highlights the presence of considerable individual sample variability, with the models demonstrating inconsistent performance across different subsets of the dataset.

### v3

@@ 2 Related Work
Recent efforts to incorporate external knowledge into datasets, like INFOSEEK [4] and Encyclopedic-VQA [24], consistently reveal the high cost and limited scalability of human annotation. While automation enables scale, these projects show significant human oversight remains crucial for data reliability. For instance, INFOSEEK relied on semi-automation due to cost, and Encyclopedic-VQA’s automated output still required extensive human validation. This underscores the enduring need for labor-intensive quality control, especially for complex or nuanced information.
@@ 3 The ReasonVQA Dataset
The ReasonVQA dataset is created to assess the capabilities of VQA models to answer multi-hop questions and to utilize external knowledge. It contains a large number of questions categorized into three levels of complexity: 1-hop, 2-hop, and 3-hop. To answer these questions, models must possess a robust ability to reason across multiple external facts. The overview of our framework, illustrated in Figure 2, to create ReasonVQA consists of three steps: (1) External Knowledge Integration, (2) Question Generation, (3) Dataset Construction. In the External Knowledge Integration step (Sect.3.1) we “link” annotated objects in the images to corresponding entities in external knowledge graphs. In the Question Generation step, multiple-hop questions are generated semi-automatically by traveling through the knowledge graph (Sect.3.2). In the Dataset Construction step (Sect.3.3), the question distribution is balanced to reduce bias, and all the images and questions are split into train set and test set with a similar answer distribution. The detailed workflow are described as algorithms in the supplementary material.
@@ 3.2.1 Template-based Question Generation
Figure 3 shows the overview of our question generation process. We generate questions through a template-based approach, utilizing predefined templates with fillable placeholders. This method supports the creation of low-cost, scalable systems by enabling a single template to yield many questions. Although the templates are manually crafted, they are designed to be generic enough to provide extensive coverage. To boost linguistic diversity, we also develop linguistic variations for each template. A key advantage of our proposed method is its ability to combine multiple templates into a single question, further enhancing system scalability. Furthermore, as we construct potential templates based on knowledge gleaned from the knowledge base (KB), the diversity of the templates aligns with the breadth of knowledge available in the KB.
One-hop Question Generation. Let us assume we are constructing questions around an object in an image, referred to as the main object. This main object corresponds to a Wikidata entity, from which we extract factual knowledge to generate questions. After obtaining external facts from Wikidata for the main object as described in Section 3.1, we use a predefined template for each property (or relation, interchangeably) occuring in these facts. Each template contains a placeholder, which will be filled later by the class name of the main object, resulting in 1-hop question. For Wikidata, the class name of an object is simply the value of either the “instance of” or “subclass of” property. The Figure 4 shows an example of 1-hop template question. In this case, the template is “Who designed __?” and the class name “skyscraper” is used to fill in the placeholder, producing the complete question “Who designed this skyscraper?”. Note that during the process of retrieving facts about the main object from Wikidata, we also obtain the answer to the question. For example, the answer for the mentioned question is “César Pelli”.
@@ 3.2.2 False Choice Generation
To generate challenging multiple choice questions, in addition to the correct answer, we need to create non-trivial false choices. First, we assign a question into one of four categories based on its correct answer: fixed, date, number, and literal. The fixed category contains questions that have an enclosed set of choices, such as questions about gender, continent, country, and language. The date, number and literal categories contain questions about date time, numeric values, and literal values, respectively. For the questions in the fixed categories, the false choices are randomly selected from the corresponding set of choices. For the date category, the false choices are generated by randomly selecting $N$ date values within [10,+10][-10,+10] years. For the number category, the false choices are generated by randomly selecting $N$ numbers within a specific range $[m,n]$, where $m=\frac{i}{2},$ $n=max(1.5i,$ $m+2N)$ with $i$ being the correct answer. For the literal category, the false choice is generated by retrieving $N$ values that come from the same property as the correct answer. In practice, we set $N=3$, meaning that there is up to three false choices for each correct answer.
@@ 4.1 Dataset Statistics and Quality Evaluation
The latest version of ReasonVQA consists of more than 598K images and about 4.2M questions. The scalability of our dataset depends on two key factors: (1) numerous objects share common attributes; and (2) multi-hop questions are constructed by nesting multiple templates. For instance, since many facts about building share the “height” attribute, we only need to define a single template for this property during question generation. To assess state-of-the-art approaches, we derived a subset version of our dataset, ReasonVQA-U. The main objective is to create a manageable subset of our dataset to assess and compare these models efficiently. The answer distribution balancing process was applied on this subset to obtain the balanced version ReasonVQA-B.
@@
In Table 1, we compare our dataset with various VQA datasets. In terms of the number of images, ours has the highest quantity. Moreover, we have a significantly larger number of questions compared to all the others, except GQA. To assess the quality of our dataset, we conducted a user study over 1000 randomly selected question and image pairs. The study involves 20 participants, who use English as the main working and/or studying language, to evaluate the correctness of the answers and naturalness of 50 randomly selected questions each. They were asked to rate the naturalness on a four-level scale: (1) very unnatural; (2) unnatural; (3) natural; and (4) very natural. Additionally, participants also marked questions with grammatical errors. The user study showed that $96\%$ of them are correct. The primary sources of error comes from numerical values, e.g. the area of cities, and ambiguities in the height or length of certain constructions. We also observed that 2.20% of the questions were rated as “very unnatural,” 13.9% as “unnatural,” 58.1% as “natural,” and 25.8% as “very natural,” with only 2.5% containing grammatical errors. This user study indicates that more than 83% of the questions are natural to humans, demonstrating that our dataset already has a good level of naturalness but still has room for improvement. Additional dataset statistics and results of the user study can be found in the Section 2 of the supplementary material.
We categorize each question into one or more domains based on the property used to generate the question. For instance, a question about the birthday of a famous person, derived from the property “date of birth,” can be classified under both the “person” and “history-events” domains. This enables the system to receive input as one or multiple specified domains and generate questions accordingly. We predefined 20 domains covering a wide range of human knowledge. During question generation, we select only properties associated with the queried domains, filtering out others to optimize efficiency. Figure 7 visualizes the domain distribution. The complete list of domains is available in Section 2 of the supplementary material.
@@ 4.2 Baseline Experiments and Analysis
Fine-tuning and Performance Gains. To assess the effectiveness of fine-tuning, we compared the performance of three vision-language models of varying scales before and after training on our dataset. Models were fine-tuned using direct image, question, and answer inputs. This experiment evaluates how well models can adapt to the dataset and improve their understanding of knowledge-based visual question answering. By analyzing performance gains, we demonstrate the potential of fine-tuning in enhancing model capabilities beyond their zero-shot performance.
Impact of Dataset Scale on Model Performance. To assess the performance and scalability of LLMs, we also benchmark them using the ReasonVQA at multiple size increments. By gradually increasing the number of questions and images, we aim to observe how each model’s accuracy relative to the amount of data. This approach will allow us to identify any limitations in model capacity, potential bottlenecks, and the overall robustness of each model under different data volumes. The results and visualizations are presented in Section 3 of the supplementary material.
@@ 4.3 Metrics
We report the accuracy of the LLMs to assess their performance. In open-ended scenario, where the LLM outputs are in free-text format, comparing responses directly with the ground truth answers can be challenging. To address this, we tested several string-matching methods to compare the output with the ground truth answer. We employ three approaches - exact match, substring, and semantic similarity - which offer different strategies for aligning and scoring the outputs based on varying degrees of match tolerance and answer format flexibility. The exact match score is 1 if the answer provided by a model precisely matches the ground truth, and 0 otherwise. However, since foundation models often provide free-text answers, exact match may not accurately reflect performance. The substring score is defined as 1 if the answer provided by the model is contained in the list of words from the ground truth answer, and 0 otherwise. Nevertheless, this approach falls short in many cases; for example, when the ground truth is “male” and the prediction is “a man,” the substring score is 0, which is incorrect. Therefore, the semantic similarity score is necessary to capture the semantic quality of the generated answer and provide a more meaningful evaluation of the LLMs. The score is based on the similarity of the generated answer and the ground truth computed using a widely recognized similarity all-MiniLM-L6-v2 [28], which is a compact, efficient model for semantic textual similarity tasks. The reported scores are the average values over the entire dataset. Section 3 of the supplementary material provides a more comprehensive report of the inference results.
@@ 4.4 Experimental Results
For recent models, we also reported in Figure 6 the accuracies in four string-matching approaches across different categories: the number of hop in the questions (1-hop, 2-hop, 3-hop); the availability of scene graph and the image source. Our results revealed that exact matching yields low overall accuracies because these models treat questions as open-ended, generating varied, free-text responses that, while often not entirely incorrect, lack the specificity required for exact matches (e.g., “very high” instead of “390 meters”). The substring approach, which considers all words from both the predicted answers and the ground truth, yields more positive scores. Semantic similarity matching consistently achieves higher scores, as it considers synonymous expressions, offering a more adaptable evaluation that closely aligns with human judgment.
@@ 5 Conclusions
We have introduced a novel, cost-effective, and scalable VQA dataset that requires external knowledge for answering questions, demanding minimal manual effort. Our construction framework leverages existing CV dataset annotations and external knowledge base Wikidata. While currently built from specific image sources, the framework is generic, allowing expansion to integrate images from other CV datasets and maintain question diversity. Users can also tailor the dataset to specific interests by selecting from 20 predefined knowledge domains. ReasonVQA features complex questions derived from scene and multi-hop knowledge graphs, proving highly challenging for state-of-the-art VQA models. We provide source code for easy integration of new images and knowledge sources, enabling researchers to expand and customize their own datasets. We believe this resource will be valuable for evaluating existing models and advancing more robust VQA systems.
Acknowledgements
@@ 1.1 Template-based Question Generation
[…]
@@ 12     $t\leftarrow\mathcal{T}_{s}[e_{j}]$
14 return $t\cup\penalty 10000\ \textnormal{{Generate(}}\textnormal{\emph{$e_{j+1}$}}\textnormal{{)}}$
@@ 19  $\{e_{j}\}$ = set of edges from $v$ to $\mathcal{C}_{i}$
20 $\mathcal{D}_{1}$ = $\mathcal{D}_{1}$ $\cup$ Generate($e_{0}$)
@@ 2.2 Dataset Domains
As detailed in the main paper, we categorized questions into 20 domains, outlined as follows. Figure 7 visualizes the domain distribution.
@@ 2.3 Question Evaluation by User Study
To evaluate the quality of generated questions, we conducted a user study with 1,000 randomly selected question and image pairs. Twenty participants, all proficient in English as their primary language for work or study, assessed the correctness of the answers and the naturalness of 50 randomly chosen questions each. For the naturalness, they rated the questions on a four-level scale: (1) very unnatural, (2) unnatural, (3) natural, and (4) very natural. Additionally, they were also asked to mark any questions with grammatical errors. The results indicated that 96% of selected answers are correct, 2.2% of the questions were rated as “very unnatural,” 13.9% as “unnatural,” 58.1% as “natural,” and 25.8% as “very natural,”, with only 2.5% of the questions had grammatical errors. The results are shown in Figure 8.
@@ Appendix 3 Experiments
In this section, we provide experimental details regarding different dataset sizes. We also present the average accuracies across domains, along with additional metrics for the benchmarked models, including the Standard Deviation (SD) and the Standard Error of the Mean (SEM). These scores are further analyzed across various aspects of our dataset.
3.1 Benchmark Results Across Dataset Sizes
To understand how models perform with varying dataset sizes, we conducted experiments on seven models Mantis-SigLIP [14], Mantis-Idefics2 [14], mPLUG-Owl3 [37], GPT-4o [13], LLaVA-OV [18], Qwen2.5-VL [2], PaliGemma-2-Mix [26] across a range of dataset sizes. In our experiments, we aimed to investigate how model performance varies with dataset size while controlling for sample difficulty. To this end, we first divided the full dataset into three pools based on the number of hops. For each target dataset size (e.g., $10k$, $20k$, …, $80k$ samples), we randomly sampled examples from each pool proportionally, maintaining the same ratio of as in the full dataset. This strategy ensures that each subset reflects the overall difficulty distribution, enabling us to evaluate how model performance scales with dataset size without bias toward easier or harder samples. Figure 9 illustrates the performance of various models across different dataset sizes. We observe that as the dataset grows, model performance initially improves but then declines with varying degrees of intensity. This decline likely occurs because as new samples are introduced, they bring additional domain knowledge and more complex questions, which may challenge the ability of models to maintain efficiency and accuracy under the increased data load. This observation suggests that integrating a wider variety of image sources and knowledge domains could create a more demanding benchmark, providing a more comprehensive assessment of model robustness.
Figure 9: Performance of models across different dataset sizes in open-ended (solid lines) and multiple choice (dashed lines) scenarios, with accuracy reported.
3.2 Benchmark Results Across Dataset Aspects
For a fine-grained analysis of our dataset, we evaluated the accuracy of 13 large language models (LLMs) (BLIP-2 [19], InstructBLIP [6], mPLUG-Owl2 [38], Idefics2 [17], Mantis-SigLIP [14], Mantis-Idefics2 [14], mPLUG-Owl3 [37], GPT-4o [13], LLaVA-OV [18], Qwen2.5-VL [2], PaliGemma-2 [32], PaliGemma-2-Mix [26], SmolVLM-Instruct [22]) across all domains, as shown in Figure 10. We observed that Health & Medicine proved to be the most challenging domain, achieving an average accuracy of only 20.7%. This was closely followed by Art & Creative Expressions, with an average accuracy of 22.5%. The most difficult question types consistently involved numerical values, such as inquiries about the area of a city. Furthermore, we computed Standard Error of the Mean (SEM) and Standard Deviation (SD) scores to better illustrate the variability within our dataset, as presented in Table 3 and 4. Overall, the SD values are relatively large, indicating a high degree of variability in the model performances across different aspects of our dataset. This variability indicates that ReasonVQA poses significant challenges to the models, as large SD values typically reflect a high sensitivity to the diverse types of samples or tasks within the dataset. At the same time, the relatively low SEM values (ranging from 0.1% to 0.7%) suggest that the mean accuracy scores for each model are estimated with high precision, despite the substantial variability indicated by the SD. This highlights the presence of considerable individual sample variability, with the models demonstrating inconsistent performance across different subsets of the dataset.
Figure 10: Average accuracy of models across domains.